Calculate no. of molecules and no. of atoms present in 5.60 L of ozone at N.T.P.
1
Expert's answer
2017-06-02T12:21:10-0400
Number of molecules in 1 mole is 6.022 · 10^23.
n = V/22.4
n (O3) = 5.60 / 22.4 = 0.25 mol
Number of molecules = 0.25 · 6.022 · 10^23 = 1.51 · 10^23
Number of O3 atoms = 1.51 · 10^23 · 3 = 4.53 · 10^23